Two sinusoidal waves of the same frequency travel in the same direction along a string. If ym1 = 3.9 cm, ym2 = 6.9 cm, ?1 = 0, and ?2 = ?/6 rad, what is the amplitude of the resultant wave? I got this incorrect.

Chapter 16, Proble 035 Yuur answer is partially cu  
Answer #1

y1 = 3.9 cm,
y2 = 6.9 cm,
w1 = 0, and
w2 = pi/6 rad

=> Y = (y1^2 + y2^2 +2*y1*y2*cos X )^0.5

=> Y = (3.9^2 +6.9 ^2 + 2*3.9*6.9*cos(pi/6))^0.5

= 10.46 cm

check the calculation be4 submission